OpenStudy (anonymous):

College Algebra: reviewing for Final Exam! If f(x)=x^2-9 and g(x)=x-3, then (f/g)(5)= ?

OpenStudy (anonymous):

\[f(x)=x ^{2}-9\] and \[g(x)=x-3\] then \[(\frac{ f }{ g })(5)\]

OpenStudy (solomonzelman):

that was a record glitch just now for me.

OpenStudy (solomonzelman):

Anyways, when you say, \(\Large\color{black}{ \frac{\color{red}{f}}{\color{blue}{g}}(x)}\), you are basically saying, \(\Large\color{black}{ \frac{\color{red}{x^2-9}}{\color{blue}{x-3}}(x)}\)

OpenStudy (anonymous):

Yes

OpenStudy (solomonzelman):

you can divide by (x-3) on top and bottom, (knowing how to factor the \(\large\color{black}{ x^2+9}\). and then plug in 5 for x.

OpenStudy (solomonzelman):

I mean -9, not +9. ***

OpenStudy (solomonzelman):

\(\Large\color{black}{ \frac{\color{red}{f}}{\color{blue}{g}}(x)=\frac{\color{red}{x^2-9}}{\color{blue}{x-3}}}\)

OpenStudy (anonymous):

 Like this correct?